Barney decides to quit his job as a corporate accountant, which pays $18,000 a month, and goes into business for himself as a certified public accountant.

He runs his business from his converted garage apartment, which he could rent out for $325 a month if he wasn’t using it as a home office. He must purchase office supplies worth $80 a month, and his monthly electricity bill has increased by $30 now that he is working out of his home office.

After six months of working from home, Barney has earned an average of $15,000 per month.

Instructions: Enter your answers as a whole number.

a. What are Barney’s monthly explicit costs?

     $

b. What are Barney’s monthly implicit costs?

     $

c. What are Barney’s monthly economic costs?

     $
Answer #1

a. Explicit cost = 80 + 30= 110

b. Implicit cost = 18000 + 325= 18325

c. Economic cost= 18325 + 110= 18435